OpenStudy (anonymous):

Which graph shows a line with slope m that passes through point m = 1 and P = (1,3)

OpenStudy (anonymous):

y-3 = x-1

OpenStudy (anonymous):

y = x + 2

OpenStudy (anonymous):

Is that good?

OpenStudy (anonymous):

this is the graph

OpenStudy (anonymous):

B. You can plot values of y for various values of x in that above equation and realize it will look like graph B.

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

Good.